OpenStudy (anonymous):

Write an equation of a line perpendicular to y = 1 over 4x + 4 in slope-intercept form that passes through the point (−2, 6).

hartnn (hartnn):

can you first find the slope of line 'y = 1 over 4x + 4 ' ?

hartnn (hartnn):

compare \(y = (1/4)x +4 \) with \(y= mx+c\) then you'll get slope = m

OpenStudy (the_fizicx99):

In \(\ \sf y = \color{blue}{m}x + c, \color{blue}{m}\) is the slope so in \(\ \sf y = \color{blue}{\dfrac{1}{4}}x + 4 \) what is the slope?

hartnn (hartnn):

 now can you tell, what slope = m = ... ?